OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f o g)(-5) A. 23 B. 3 C. -53 D. -9

OpenStudy (anonymous):

@DJ♬Øϟϟ¥

OpenStudy (anonymous):

(f o g)(-5) = f(g(-5) can you take it on from there?

OpenStudy (anonymous):

No i have no idea

OpenStudy (anonymous):

Yes So first find g(-5) by replacing x by -5 in g(x) So we get g(-5)= -4(-5) + 3 = 20 + 3 = 23 So g(-5)= 23 Since we know above formula f(g(-5))= f(23)= -2(23) - 7 = -53

OpenStudy (anonymous):

okay thank you ! ^_^